if the pair of lines 2x+3y+7=0 and ax+by+14=0 are coincident lines then what is the value of ‘a’ and ‘b’  are respectively equal to
 

Grade:10

1 Answers

Shivam shukla
2 years ago
There are many ways of doing this que 
As we can see in the que. if we multiply given equation by 2 we see the constant coefficients of both the equation became equal. And hence coefficients of x and y will also be equal for both the lines to be coincident.
i.e. a=4.   &     b=6.
2nd way
First of all we find two points from the given equations at x is equal to zero and at Y is equal to zero.
i.e. at x=0
2*0+3y=-7
Y= -7/3.  (0,-7/3)
N at y=0
2x=-7
X=-7/2.   i.e. (-7/2,0)
further these points will satisfy the given second equation of coincident line
And we'll get values of a and b .
i.e.
At (0,-7/3)
-7b/3 = -14.       b = 6
At (-7/2,0)
-7a/2= -14.         a = 4
 
There are many more ways of doing the question. 
Thank you